### v1

@@ 2.4 Equilibrium: Tariff Pass-Through Elasticities
Our pricing framework provides a basis for analyzing the short-run impact of an unanticipated tariff imposed on imports from country $c$. Let $T_{c}$ denote the gross tariff rate. The (log) price that exporter $i$ from country $c$ charges importer $j$ is given by:
$\ln p_{ij}=\ln\mu_{ij}+\ln c_{i}+\ln T_{c}.$
While a tariff on country $c$ may, in principle, affect all trade relationships involving exporters from $c$, we focus on the short-run, direct effect at the relationship level. We interpret the tariff shock as a small, unanticipated perturbation around the initial equilibrium. Accordingly, we treat $T_{c}$ as a pair-specific shock and hold constant prices in all other relevant matches such that $dp_{rj}=0$ for all $r\in\mathcal{Z}_{j}\setminus{i}$ and $dp_{i\ell}=0$ for all $\ell\in\mathcal{Z}_{i}\setminus{j}$.
@@
$\Phi_{ij}\equiv\frac{d\ln p_{ij}}{d\ln T_{c}}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$ (2.11)
@@
$\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}\quad\textit{and}\quad\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$
denote the partial elasticities of the equilibrium markup $\mu_{ij}$ and the exporter’s marginal cost $c_{i}$ to changes in the bilateral price $p_{ij}$, respectively.
@@
Proposition 3 highlights two distinct mechanisms shaping tariff pass-through in firm-to-firm trade: a markup channel, capturing strategic pricing responses, and a cost channel, capturing how the exporter’s marginal cost adjusts in response to price changes. In models with CES demand, monopolistic competition, and constant marginal costs, both $\Gamma_{ij}$ and $\Lambda_{ij}$ are zero, yielding full pass-through ($\Phi_{ij}=1$). More generally, $\Phi_{ij}$ may lie above, below, or at one depending on the magnitudes of $\Gamma_{ij}$ and $\Lambda_{ij}$.
2.4.1 Markup Elasticity
The markup elasticity $\Gamma_{ij}$ measures how the equilibrium markup $\mu_{ij}$ responds to changes in the bilateral price $p_{ij}$. It is given by:
[…]
where $\omega_{ij}^{\Gamma}\equiv\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\in[0,1]$, $\Gamma_{ij}^{\text{oligopoly}}$ denotes the oligopoly markup elasticity, $\Gamma_{ij}^{\text{oligopsony}}$ the oligopsony markup elasticity, and $\Gamma_{ij}^{\omega}$ the elasticity of the bargaining weight with respect to the bilateral price.
The term in square brackets captures the direct elasticity of the markup, expressed as a convex combination of the oligopoly and oligopsony components. The second term reflects how the bargaining weight $\omega_{ij}$ itself responds to price changes. We discuss each component in turn.
Oligopoly Markup Elasticity
The oligopoly markup elasticity dominates when $\phi\to 0$, in which case $\Gamma_{ij}\to\Gamma_{ij}^{\text{oligopoly}}$. It is given by
$\Gamma_{ij}^{\text{oligopoly}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln p_{ij}}=\frac{1}{\varepsilon_{ij}-1}\cdot\frac{\rho-\varepsilon_{ij}}{\varepsilon_{ij}}\cdot(\rho-1)(1-s_{ij})\geq 0.$
This elasticity reflects the standard logic of strategic complementarities in price-setting: when a tariff raises the bilateral price $p_{ij}$, exporter $i$ reduces its markup to limit trade diversion, resulting in incomplete pass-through (see Amiti et al., 2014; Auer and Schoenle, 2016; Garetto, 2016; Amiti et al., 2019a). The response is strongest at intermediate supplier shares, giving rise to a U-shaped relationship between $\Gamma_{ij}^{\text{oligopoly}}$ and $s_{ij}$. [^12] Amiti et al. (2014) shows that, to a first-order approximation, pass-through decreases with the exporter’s share through strategic complementarities. However, we cannot rely on the same approximation, given our focus on bilateral markets, where both very low and very high market shares are observed in the data.
Oligopsony Markdown Elasticity
The oligopsony markdown elasticity dominates when $\phi\to 1$, in which case $\Gamma_{ij}\to\Gamma_{ij}^{\text{oligopsony}}$. It is given by
@@ 2.4.1 Markup Elasticity
$\Gamma_{ij}^{\text{oligopsony}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln p_{ij}}=\left(\frac{x_{ij}\cdot(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\cdot(1-(1-x_{ij})^{\frac{1}{\theta}})}-1\right)(1-x_{ij})\,\varepsilon_{ij}\leq 0.$
@@
This elasticity captures the logic of strategic substitutabilities among importers. [^13] Strategic substitutabilities arise because a decline in demand by other buyers lowers marginal costs and prices, encouraging importer $j$ to expand purchases. As the price $p_{ij}$ rises, importer $j$ reduces demand, shrinking the exporter’s quasi-rents. This weakens the importer’s bargaining position, reducing the markdown and amplifying the price response, potentially resulting in more-than-complete pass-through.
@@
The elasticity $\Gamma_{ij}^{\text{oligopsony}}$ depends on both the importer’s buyer share $x_{ij}$ and the exporter’s supplier share $s_{ij}$. It is U-shaped in $x_{ij}$: the elasticity vanishes when $x_{ij}\to 0$ (atomistic buyer) or $x_{ij}\to 1$ (monopsonist), and reaches its peak at intermediate values. It also declines with $s_{ij}$, since a higher supplier share reduces the demand elasticity $\varepsilon_{ij}$, limiting changes in $x_{ij}$ and weakening the markdown response.
The Role of the Endogenous Bargaining Weight
The elasticity $\Gamma_{ij}^{\omega}\equiv\frac{d\ln\omega_{ij}}{d\ln p_{ij}}$ captures how the bargaining weight $\omega_{ij}$ responds to price changes. Its sign and magnitude depend on the exporter’s supplier share $s_{ij}$ and are derived in Appendix A.3. For empirically relevant values of $s_{ij}$ and parameter ranges, this elasticity is typically small. In particular, as $s_{ij}$ approaches zero or one, where $\omega_{ij}$ converges to the importer’s exogenous bargaining strength $\phi$, the elasticity $\Gamma_{ij}^{\omega}$ goes to zero.
2.4.2 Cost Elasticity
The cost elasticity $\Lambda_{ij}$ measures how exporter $i$’s marginal cost $c_{i}$ responds to changes in the bilateral price $p_{ij}$, via adjustments in traded quantity. While prior work emphasizes the role of the cost channel in shaping pass-through (e.g., Burstein and Gopinath, 2015; Amiti et al., 2019a), our contribution is to show that this elasticity varies systematically across matches through both demand and supply channels.
Formally:
$\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ $=\frac{d\ln c_{i}}{d\ln q_{ij}}\cdot\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
$=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}\geq 0.$
The elasticity declines with the exporter’s supplier share $s_{ij}$, which reduces the residual demand elasticity $\varepsilon_{ij}$, and increases with the importer’s buyer share $x_{ij}$, which amplifies the residual supply elasticity $\frac{d\ln c_{i}}{d\ln q_{ij}}$. Importantly, this mechanism operates independently of bargaining power ($\phi$), and instead reflects the concentrated nature of the trade network.
2.4.3 Pass-Through Elasticity and Bilateral Market Shares
@@
Notes: The figure presents heatmaps of the pass-through elasticity $\Phi_{ij}$ across combinations of $s_{ij}$ (x-axis) and $x_{ij}$ (y-axis), under alternative assumptions about $\phi$ and $\theta$. We set $\theta=0.5$ in the decreasing returns case and $\phi=0.5$ in the intermediate bargaining case. Other parameters are fixed at $\gamma=0.5$, $\varrho=1$, $\nu=4$, and $\rho=10$.
The interaction between markup and cost channels generates rich heterogeneity in pass-through elasticities across matches, shaped by bargaining power ($\phi$), returns to scale ($\theta$), and bilateral market shares ($s_{ij}$ and $x_{ij}$). Figure 2 illustrates these patterns using heatmaps of the pass-through elasticity $\Phi_{ij}$ as a function of supplier share ($s_{ij}$) and buyer share ($x_{ij}$), under three bargaining regimes—$\phi\to 0$, $\phi\in(0,1)$, and $\phi\to 1$—with decreasing returns to scale ($\theta<1$) in the top row and constant returns to scale ($\theta=1$) in the bottom row.
Several key insights emerge. First, pass-through $\Phi_{ij}$ increases with importer bargaining power $\phi$. As $\phi$ rises, the markup elasticity shifts toward the oligopsony markdown elasticity, increasing pass-through through strategic substitutabilities. This is most visible in the shift from the left to the right columns within each row.
Second, when $\theta<1$, pass-through is incomplete across most of the $(s_{ij},x_{ij})$ space. This is because the cost elasticity $\Lambda_{ij}$, which is always non-negative, dominates the overall pass-through elasticity $(\Gamma_{ij}+\Lambda_{ij})$ when $\Gamma_{ij}$ is negative due to strong buyer power. Moreover, pass-through values under $\phi\in(0,1)$ and $\phi=1$ are nearly indistinguishable, indicating that pass-through is relatively insensitive to the exact level of $\phi$ in this regime. This reflects two forces: (i) the cost channel is strong when returns to scale are decreasing and importers’ bargaining power is high, and (ii) the markup elasticity $\Gamma_{ij}$ is either muted by offsetting effects of strategic complementarities and substitutabilities or generally low in values. [^14] For more details, see Appendix A.4.
Third, pass-through $\Phi_{ij}$ depends on $x_{ij}$ only when $\theta<1$, in which case it declines with $x_{ij}$. This yields a sharp empirical prediction: a negative relationship between pass-through and buyer share $x_{ij}$ implies decreasing returns to scale, regardless of the level of bargaining power. This result is formalized below.
@@
If pass-through $\Phi_{ij}$ decreases with the importer’s buyer share $x_{ij}$, then $\theta<1$.
Proof: See Appendix A.4.
By contrast, the relationship between $\Phi_{ij}$ and the supplier share $s_{ij}$ is less clear-cut. As shown in Figure 2, pass-through declines with $s_{ij}$ primarily when $\phi$ is low and $\theta$ is near one, a setting in which the markup channel dominates and cost-based adjustments are limited. In this region, strategic complementarities give rise to a U-shaped relationship between $\Phi_{ij}$ and $s_{ij}$. As either $\phi$ or $x_{ij}$ increases, the cost channel becomes more prominent, and pass-through tends to rise with $s_{ij}$. While we do not formally characterize the conditions under which this reversal occurs, we show below that $\phi$ is weakly increasing in $s_{ij}$. This monotonicity supports interpreting the empirical evidence as inconsistent with the low-$\phi$, high-$\theta$ case.
@@ 2.5 Discussion
This section reviews key modeling assumptions and discusses potential extensions.
2.5.1 Bargaining Protocol and Quantities
The baseline model assumes demand-determined quantities, whereby the importer chooses input quantities to minimize total cost, taking the price as given. This assumption provides analytical tractability, yields closed-form solutions for markups and pass-through, and nests standard models of international trade. An important implication is that prices are allocative, a feature that aligns with our empirical findings and with firm-level evidence in related settings (Gopinath and Itskhoki, 2011).
We consider two alternative bargaining protocols as benchmarks for understanding the role of quantity determination. The first is efficient bargaining, discussed more formally in Appendix B.1. This case corresponds to the vertically integrated benchmark, where importer and exporter jointly negotiate over both price and quantity to maximize total surplus. While theoretically appealing, this setup implies that prices are non-allocative transfers, inconsistent with the empirical evidence. Moreover, the vertically integrated case may be unrealistic in the context of arm’s-length firm-to-firm trade, where limited commitment and contracting frictions are prevalent (Antràs, 2020).
The second alternative, detailed in Appendix B.2, is supply-driven bargaining: the exporter sets quantity for a given price, and price is then negotiated based on the resulting supply curve. As $\phi\to 1$, this nests the classic monopsony benchmark commonly used in labor markets (e.g., Berger et al., 2022). The key distinction between this setup and our baseline lies in their welfare implications: in the supply-driven case, buyer power lowers both prices and quantities, while supplier power helps restore efficiency. In contrast, under demand-driven bargaining, buyer power mitigates upstream distortions and improves efficiency. [^15] See Avignon et al. (2024) and Demirer and Rubens (2025) for recent discussions of how quantity-setting assumptions affect welfare outcomes.
While both alternatives offer useful benchmarks, they are less suited to the goals of this paper. The supply-driven model, in particular, does not provide closed-form expressions for key variables, which makes it challenging to generate clear predictions or link the model to data compared to our baseline setup. Still, it remains a valuable direction for future work, especially for studying welfare effects when buyer power leads to inefficiencies.
2.5.2 Outside Options
In our model, each firm’s outside option reflects payoffs from trade with all other existing partners, excluding the focal match. This assumption allows us to express markups and pass-through in terms of market shares and a small set of parameters, facilitating structural estimation and counterfactual analysis.
While analytically convenient, the assumption may appear restrictive if disagreement leads a firm to form or sever other relationships. Appendix B.3 explores a more flexible setup in which disagreement affects the importer’s cost and the exporter’s revenue non-parametrically. Although more realistic, this extension introduces an identification problem: the parameters $\phi$ and $\theta$ can no longer be separately identified from the outside option, which limits their interpretability and empirical tractability.
2.5.3 General Equilibrium Forces
We conclude the theoretical section by noting that all results are derived under a partial equilibrium approach, holding fixed general equilibrium variables such as wages, demand shifters, and competitor behavior. This approach simplifies the analysis but also raises questions about the role of general equilibrium adjustments in the results.
General equilibrium considerations are critical for interpreting pass-through elasticities. The structural pass-through elasticity derived in Section 2.4 is a direct or partial elasticity, holding constant aggregate variables such as wages, demand conditions, and, importantly, competitors’ prices and sourcing decisions. In reality, tariff shocks may also affect variables such as foreign wages or domestic export prices. If these effects are not fully observed or controlled for, reduced-form estimates may conflate the direct impact with general equilibrium responses, making it harder to align empirical pass-through coefficients with model-based elasticities (Burstein and Gopinath, 2015). [^16] In Appendix B.4, we extend equation (2.11) to account for indirect effects, such as how a shock to exporter $i$ influences other prices and quantities, which may in turn affect $p_{ij}$. While we do not model full general equilibrium dynamics, this extension illustrates how spillovers across relationships may lead to reduced-form estimates diverging from structural ones. Therefore, our goal is not to recover structural elasticities from reduced-form regressions, which would be inappropriate in this context (Berger et al., 2022).
Instead, we pursue a structural approach, which helps mitigate similar concerns. Our approach has three advantages. First, the theoretical relationship between bilateral shares, markups, and pass-through holds parametrically regardless of the specific general equilibrium environment. While the parameter values depend on how aggregate variables such as tariffs, demand shifters, and wages co-move in equilibrium, we do not need to model their relationships explicitly. Instead, identification is based on the cross-sectional variation of prices across importers within exporter–product–year cells, without making assumptions on the general equilibrium environment.
Second, we can also isolate and test the cross-sectional predictions of the model independently of general equilibrium forces. In particular, even if aggregate variables shift over time, the model predicts specific patterns between firm-to-firm shares, markups, and pass-through at a given point in time. These predictions can be assessed empirically using within-period variation across matches, which we exploit in the next sections.
Third, we can evaluate the performance of the estimated model by comparing its ability to replicate observed price changes in response to tariff shocks, thereby gauging the significance of general equilibrium forces. As shown in the empirical analysis, the estimated model fits the observed price changes well. This suggests that the short-run effects of tariffs on prices can be understood primarily through the lens of partial equilibrium mechanisms, and supports the usefulness of our framework for studying firm-to-firm pricing and pass-through.

### v2

@@ 2.4 Pass-Through Characterization
Our pricing framework allows for an analytical characterization of the short-run effects of unanticipated tariffs. Let $T_{c}$ denote the gross tariff rate. We assume that the tariff enters multiplicatively into the price charged by exporter $i$ from country $c$ to importer $j$:
$\ln p_{ij}=\ln\mu_{ij}+\ln MC_{i}+\ln T_{c(i)}.$
While a tariff on country $c$ may affect all relationships involving its exporters, we treat it as a pair-specific shock by holding all other match prices fixed: $dp_{rj}=0$ for all $r\in\mathcal{Z}_{j}\setminus{i}$ and $dp_{i\ell}=0$ for all $\ell\in\mathcal{Z}_{i}\setminus{j}$.
@@
$\Phi_{ij}\equiv\frac{d\ln p_{ij}}{d\ln T_{c(i)}}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$ (2.15)
@@
$\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}\quad\textit{and}\quad\Lambda_{ij}\equiv-\frac{d\ln MC_{i}}{d\ln p_{ij}}$
denote the partial elasticities of the equilibrium markup $\mu_{ij}$ and the exporter’s marginal cost $MC_{i}$ to changes in the bilateral price $p_{ij}$, respectively.
@@
Proposition 3 highlights two distinct mechanisms shaping tariff pass-through in firm-to-firm trade: a markup channel, capturing strategic markup responses, and a cost channel, capturing how the exporter’s marginal cost adjusts in response to price changes. [^14] Appendix B.4 extends equation (2.15) to incorporate indirect effects, for example shocks to exporter i that affect other prices and quantities and in turn $p_{ij}$. While we do not model these dynamics explicitly, the extension illustrates how spillovers across relationships can drive a wedge between reduced form and structural estimates. We characterize each channel in turn.
The Markup Channel ($\Gamma_{ij}$).
The elasticity $\Gamma_{ij}$ captures strategic markup responses to price changes. It can be written as
@@
$\Gamma_{ij}=\left[(1-\omega_{ij}^{\Gamma})\Gamma_{ij}^{\text{oligopoly}}+\omega_{ij}^{\Gamma}\Gamma_{ij}^{\text{oligopsony}}\right]+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega},$
where $\omega_{ij}^{\Gamma}\equiv\omega_{ij}\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\in[0,1]$, $\Gamma_{ij}^{\text{oligopoly}}$ and $\Gamma_{ij}^{\text{oligopsony}}$ denote the oligopoly and oligopsony markup elasticities, and $\Gamma_{ij}^{\omega}$ is the elasticity of the bargaining weight with respect to the bilateral price.
The term in square brackets captures the direct markup adjustment, holding bargaining weights ($\omega_{ij})$ fixed. The second term reflects the endogenous response of these weights.
When exporter bargaining power dominates ($\phi\to 0$), $\Gamma_{ij}$ converges to $\Gamma_{ij}^{\text{oligopoly}}$, given by
$\Gamma_{ij}^{\text{oligopoly}}=-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln p_{ij}}=\frac{1}{\varepsilon_{ij}-1}\cdot\frac{\rho-\varepsilon_{ij}}{\varepsilon_{ij}}\cdot(\rho-1)(1-s_{ij})\geq 0.$
In this regime, the markup channel operates through strategic complementarities, as is standard in oligopoly models. [^15] See, e.g., amiti2014importers; auer2016market; garetto2016firms; amiti2019international. An increase in the bilateral price induces the exporter to compress its markup in order to limit trade diversion, generating incomplete pass-through. The response is strongest at intermediate supplier shares, implying a U-shaped relationship between $\Gamma_{ij}^{\text{oligopoly}}$ and $s_{ij}$.161616amiti2014importers show that, in their model, a first-order approximation around small supplier shares implies that pass-through declines with exporter share. We cannot rely on this approximation in our bilateral pricing framework, where both low and high shares are observed.
When importer bargaining power dominates ($\phi\to 1$), $\Gamma_{ij}$ converges to $\Gamma_{ij}^{\text{oligopsony}}$, given by
$\Gamma_{ij}^{\text{oligopsony}}=-\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln p_{ij}}=\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)(1-x_{ij})\varepsilon_{ij}\leq 0.$
@@
In this case, the elasticity $\Gamma_{ij}$ is governed by markdown adjustments and reflect strategic substitutabilities among importers. As $p_{ij}$ rises, importer $j$ contracts demand, reducing quasi-rents and weakening its bargaining position. The resulting markdown compression amplifies the price response and can generate more-than-complete pass-through.
@@
The elasticity $\Gamma_{ij}^{\text{oligopsony}}$ depends on both the buyer share $x_{ij}$ and the supplier share $s_{ij}$. It is U-shaped in $x_{ij}$, vanishing as $x_{ij}\to 0$ (atomistic buyer) or $x_{ij}\to 1$ (monopsonist), and peaking at intermediate values. It declines in $s_{ij}$ because a higher supplier share reduces the demand elasticity $\varepsilon_{ij}$, limiting variation in $x_{ij}$ and weakening the markdown response.
With two-sided market power $(\phi\in(0,1))$, the elasticity $\Gamma_{ij}$ reflects a convex combination of oligopoly and oligopsony markup elasticities, with weights $\omega_{ij}^{\Gamma}$ proportional to importer bargaining power. It also reflects the endogenous response of bargaining weights through $\Gamma_{ij}^{\omega}\equiv\frac{d\ln\omega_{ij}}{d\ln p_{ij}}$, whose sign and magnitude depend on $s_{ij}$. [^17] The exact expression for this term is provided in Appendix A.3. As $s_{ij}\to 0$ or $s_{ij}\to 1$, $\omega_{ij}$ converges to the exogenous parameter $\phi$, implying $\Gamma_{ij}^{\omega}\to 0$.
The Cost Channel ($\Lambda_{ij}$).
The cost elasticity $\Lambda_{ij}$ captures how marginal costs respond to price changes. An increase in prices reduces demand and, under decreasing returns, lowers marginal costs. Formally,
$\Lambda_{ij}=\frac{1-\theta}{\theta}\,x_{ij}\,\varepsilon_{ij}\geq 0.$ (2.16)
The first term, $(1-\theta)/\theta$, captures the slope of marginal cost or, equivalently, the inverse export supply elasticity at the exporter level. The buyer share $x_{ij}$ captures the exposure of the exporter’s total output to the relationship: importers that account for a greater fraction of the exporter’s output exert stronger scale effects on marginal cost. The residual demand elasticity $\varepsilon_{ij}$ captures how strongly bilateral quantity responds to a change in price. The product $(1-\theta)/\theta\cdot x_{ij}$ is the residual inverse export supply elasticity faced by buyer $j$: it scales with the buyer share, reflecting each buyer’s exposure to the exporter’s overall supply curvature. Together, $x_{ij}$ and $\varepsilon_{ij}$ determine the strength of the match-specific scale effect.
This mechanism may appear similar to a standard cost channel driven by firm-level decreasing returns. While upward-sloping export supply can generate incomplete pass-through, a model with atomistic buyers attributes this entirely to supply curvature. As shown in Appendix B.5, this implies uniform pricing and identical pass-through across buyers sourcing from the same exporter. In contrast, our framework features non-atomistic buyers and produces relationship-specific residual cost elasticities that scale with buyer shares, generating cross-sectional heterogeneity in pass-through.
Pass-Through Elasticity and Bilateral Market Shares
Figure 2 plots the pass-through elasticity $\Phi_{ij}$ as a function of $s_{ij}$ and $x_{ij}$ for three values of $\theta$ (rows) and $\phi$ (columns). [^18] We set $\theta=0.9$ as an intermediate value of the returns-to-scale parameter, consistent with estimates in the production function literature, typically between 0.9 and 1; see, e.g., basu1997returns. The figure shows that the same primitives governing oligopsony power also shape pass-through. Variation in $\theta$ (vertical) captures how exporters’ returns to scale, via the inverse export supply elasticity, determine the balance between cost and markup channels, while variation in $\phi$ (horizontal) captures how bargaining power governs strategic markup adjustment.
@@
Notes: The figure presents heatmaps of the pass-through elasticity $\Phi_{ij}$ as a function of market shares $s_{ij}$ (x-axis) and $x_{ij}$ (y-axis), for $\phi\in\{0,0.5,1\}$ (columns) and $\theta\in\{0.5,0.9,1\}$ (rows). Other parameters are fixed at $\gamma=0.5$, $\varrho=1$, $\nu=4$, and $\rho=10$.
Moving from top to bottom, $\theta$ rises from 0.5 to 1. At $\theta=0.5$, export supply is steep, so marginal cost is highly sensitive to output and the cost channel dominates across the $(s_{ij},x_{ij})$ space, rendering pass-through largely invariant to $\phi$ and the markup channel. As $\theta$ increases, the cost channel weakens and markup adjustment becomes more important. Markdown responses also attenuate, but the cost channel continues to dominate in $x_{ij}$ (Appendix A.4), so pass-through becomes less incomplete. At $\theta=1$, supply is flat, the cost channel vanishes, and pass-through is governed entirely by markup adjustment.
Moving left to right, higher $\phi$ traces the role of bargaining power in shaping strategic markup adjustment. At $\phi=0$, the markup channel reflects strategic complementarities that dampen pass-through; at $\phi=1$, it reflects strategic substitutabilities associated with oligopsony power that amplify pass-through. This gradient is most pronounced in the bottom row ($\theta=1$), where only the markup channel operates, and attenuates as $\theta$ falls. At $\theta=0.5$, markup adjustment is second order.
A key implication of Figure 2 is that pass-through is incomplete over most of the parameter space, with the sole exception of the constant-returns, full-importer-power case ($\theta=1$, $\phi=1$). Different $(\theta,\phi)$ pairs generate incomplete pass-through through distinct mechanisms and to varying degrees. In the lower-left region of the heatmaps–corresponding to the low-share range where most bilateral matches lie in the data (Table 1)–average pass-through is similar across specifications, making models difficult to distinguish on that margin alone. By contrast, the sign and magnitude of the pass-through gradient in $x_{ij}$ differ sharply across variants. The following proposition formalizes this contrast, which we exploit in Section 4.3 to discriminate among models using cross-sectional pass-through moments.
@@
When $\theta<1$, the pass-through elasticity $\Phi_{ij}$ decreases with the importer’s buyer share $x_{ij}$.
Proof: See Appendix A.4. Intuitively, when $\theta<1$, the cost channel dominates the markdown channel. Equation (2.16) implies that pass-through falls with $x_{ij}$.
The relationship between $\Phi_{ij}$ and $s_{ij}$ is, by contrast, ambiguous. Under strategic complementarities, pass-through is U-shaped in $s_{ij}$. Under strategic substitutabilities and through the cost channel, it is increasing in $s_{ij}$. The former pattern arises when bargaining power is low and returns to scale are close to constant, while the latter emerges as either force strengthens. Which effect dominates is ultimately an empirical matter.
@@ 2.5 Discussion
We briefly review our key modeling choices before turning to the empirical analysis.
Bargaining Protocol and Quantities
Our baseline model assumes demand-driven quantities, whereby the importer chooses quantities given negotiated prices. This ensures allocative pricing and nests the standard monopolistic-competition benchmark as a limiting case, consistent with firm-level evidence in international trade (e.g., gopinath2011search).
We derive two alternative protocols in the Appendix. The first is efficient bargaining (Appendix B.1), which maximizes joint surplus but implies non-allocative prices, a prediction at odds with the data. The second is supply-driven bargaining (Appendix B.2), a monopsony-style benchmark, which implies that prices and quantities move together along the exporter’s supply curve. This implication is inconsistent with our empirical evidence that tariff-induced price increases are accompanied by quantity declines.
One caveat of our framework is that when $\phi$ is sufficiently high, equilibrium prices can fall below the efficient level ($p<p^{\text{eff}}$), implying quantities above the competitive benchmark ($q>q^{\text{eff}}$). Recent theories of bilateral oligopoly rule out this case by imposing a supplier participation constraint that prevents quantities from exceeding the exporter’s supply (marginal cost) curve when $p<p^{\text{eff}}$ (e.g., demirer2025welfare; avignon2024markups). We abstract from this extension because, at high $\phi$, it places equilibrium on the upward-sloping supply curve, implying positive price-quantity comovement, contrary to our evidence.
General Equilibrium and Network Spillovers
Our theoretical results are derived in partial equilibrium, holding wages, aggregate demand shifters, and competitors’ prices and quantities fixed. The pass-through elasticity derived in Section 2.4 therefore isolates the direct, intensive-margin response of markups and marginal costs to a tariff shock.
Embedding this framework in a general equilibrium environment introduces two main complexities. First, under decreasing returns, quantity adjustments along one link affect the exporter’s marginal cost across all links. Second, aggregate shocks or network spillovers may differentially affect outside options across matches of the same exporter. Appendix B.3 discusses how generalizing outside options modifies the bargaining problem. Fully accommodating these channels, however, requires committing to a specific macroeconomic closure for wages, final demand, and network formation, which is beyond the scope of this paper. As we show below, the partial equilibrium approach is supported by the close fit between model predictions and untargeted pass-through moments.